Evaluate the expression leaving the answer as an integer or fraction
(10)-3..10 to the -3 power

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(48558)

Evaluate the expression leaving the answer as an integer or fraction
(10)-3..10 to the -3 power
----------------------------------------------------
10^(-3) = (10^3)^(-1) = 1000^(-1) = 1/1000
============================================
Cheers,
Stan H.

Answer by solver91311(12126)